OpenStudy (anonymous):

how would u solve: sq rt of 24rs times sq rt of 6r^3s^4 times sq rt of rs^2

OpenStudy (anonymous):

\[\sqrt[2]{24rs}\times \sqrt[2]{6r^3s^4}\times \sqrt[2]{rs^2}\]

OpenStudy (radar):

Sqrt(24rs x 6r^3s^4 x rs^2)

OpenStudy (radar):

\[\sqrt{144 xr ^{5}s ^{7}}\]

OpenStudy (radar):

\[12r ^{2}s ^{3}\sqrt{rs}\]

OpenStudy (anonymous):

but how did u get that?

OpenStudy (radar):

The first equartion placed all the terms in one radical becoming factors producing the second line.

OpenStudy (radar):

The third line shows taking the perfect squares and placving their roots outside the radical.

OpenStudy (radar):

leaving only the factors which were not perfect squares in the radical. Note you can start from the finish and work backwards and come up with your original expression. The 12 goes back as 144 the r^2 goes back as r^4 ......etc.

OpenStudy (anonymous):

i see. thank u